# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

### 19 Mar 2026 v7

@@ 9 Conclusion
We formalized an orientation boundary for first-order rewrite systems with step-duplicating recursors. Twelve direct measure classes are excluded by schema-level impossibility theorems (Theorems 4.1 to 4.13); the escape side is classified by the DP clause, the explicit DP base-order boundary, transparency essentiality, the generalized bounded-polynomial escape condition, the escape trichotomy, and the two SCC synchronization theorems; and the concrete witness calculus KO7 carries a guarded certification chain together with full-step polynomial and MPO witnesses.
The theory is explanatory as well as negative. The barrier stack separates baseline additive/compositional failure, widened scalar families, vector / pair projection layers, and the symbolic variable-condition side; the escape results match that organization closely. It is a boundary theorem for explicit direct families, not a claim that KO7 defeats termination methods wholesale. Dependency pairs, LPO, the nonlinear witness, and the KO7-specialized MPO remain on the successful side of the boundary, and the repository makes those imports explicit at the theorem level.
Open directions include extending the projection-based trichotomy layer to broader mixed matrix and tropical families, tightening the root-normalizer and SafeStepCtx upper envelopes from the current explicit tower and fast-growing-family bounds toward a sharper slow-growing extraction from the $\omega^{\omega}\!\cdot\!2$ calibration, replaying a meaningful fragment of the external certificate chain inside Lean, and lifting the barrier analysis to higher-order or sharing-based rewriting semantics.
@@ A.2 Module map
Table LABEL:tab:module-map lists the theorem-bearing kernel and meta modules in the active artifact. Lightweight test files such as Test/Sanity.lean are intentionally omitted.
@@
Meta/ContextClosedBarrier Root-to-context barrier survival corollaries for StepCtxFull (stepCtxFull_orientation_implies_root,
no_stepCtxFull_orientation_affine_with_pump)
@@
Meta/SafeStep_Complexity_Ordinal SafeStepCtx derivation-length bound via ctxFuel; size-indexed tower-exponential complexity bound (safestep_length_bounded_by_size)
Meta/SafeStep_Complexity_FastGrowing Explicit finite fast-growing envelope majorizing the same SafeStepCtx length bound (fgOmegaEnvelope, safestep_length_bounded_by_fgOmegaEnvelope)
Meta/SafeRoot_Complexity Exact-path realization and explicit upper envelope for the certified root normalizer (normalizeSafeSteps_realized,
normalizeSafeSteps_le_complexity_bound)
Meta/Reachability_Complexity Explicit decision-cost envelope for safe-normal-form reachability, together with a linear lower-bound family (reachabilityDecisionCost_le_complexity_bound,
reachabilityDecision_has_linear_lower_family)
Meta/SafeStepCtx_Confluence Discharged global contextual local-join obligation and unconditional SafeStepCtx confluence, retaining the exact Newman equivalence (localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll)
@@
Meta/Impossibility_Lemmas Failed-measure catalog; historical dead-end notes moved to Notes/Impossibility_DeadEnds.md
@@
Meta/CompositionalMeasure_Impossibility KO7 instantiation; KO7 DP witness; KO7 affine barrier (Thms. 4.184.21;
@@
Meta/KBO_Impossible Explicit KO7 KBO-style impossibility corollary extracted from the symbolic barrier (no_kbo_orients_ko7_rec_succ)
@@
Meta/DependencyPairs_Fragment Reusable DP fragment: rank-based pair problems and SCC-path descent (DPProjection.wfRev, SCCCycle.target_lt_source)
@@
Meta/PolyInterpretation_FullStep Nonlinear polynomial $W$; axiom-violation witnesses (Rem. 4.20;
@@ A.3 Relation/property status matrix
Step Root termination: yes Root confluence: no (witness); normalizer: n/a nonlinear polynomial and MPO witnesses; MPO ranking bounded below $\varphi_{7}(0)$ wf_StepRev_poly, wf_StepRev_mpo, wf_StepRev_mpo_below_veblen7, not_localJoinStep_eqW_refl
SafeStep Strong normalization: yes Root confluence + certified normalizer + reachability to safe normal-form targets triple-lex ordinal upper-bound calibration; explicit root-normalizer lower/upper envelope wf_SafeStepRev_c, confluentSafe, normalizeSafe_sound, reachability_decidable, normalizeSafeSteps_le_complexity_bound, safeMeasure_below_epsilon0
SafeStepCtx Partial-context SN: yes Confluence: yes; exact Newman equivalence retained size-indexed tower bound plus explicit fast-growing-family envelope wf_SafeStepCtxRev, localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, safestep_length_bounded_by_size, safestep_length_bounded_by_fgOmegaEnvelope
@@ A.4 Claim-to-code index
Tables LABEL:tab:claim-code-barrier and LABEL:tab:claim-code-certification index the numbered paper results by Lean identifier, abstraction level, and side conditions. This complements the module map: the tables below are organized by paper statement rather than by file.
@@
Thm. 4.7 no_max_orients_dup_step_of_unbounded, no_global_step_orientation_max_of_succ_pump Schema / KO7 Max-plus constructor-local interface with unbounded range, or positive successor / left-wrapper drift.
Cor. 4.8 no_arctic_primary_orients_dup_step_of_unbounded, no_global_step_orientation_arctic_primary_of_unbounded Schema / KO7 Distinguished finite arctic primary projection satisfying the max-plus interface.
Thm. 4.9 no_matrixD_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with one tracked affine coordinate and growth in that coordinate.
Thm. 4.10 no_matrix2_lex_orients_dup_step_of_unbounded_primary Schema / KO7 Lexicographic order by the first component; unbounded or pumped primary coordinate.
Thm. 4.11 no_matrixMix2_orients_dup_step_of_sum_pump Schema / KO7 Balanced mixed $2\times 2$ constructor maps and an unbounded aggregate-sum pump.
Thm. 4.12 no_matrixFunctional_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with a fixed nonzero natural-weighted scalar projection satisfying the affine interface.
Thm. 4.13 no_orients_dup_step_of_scalar_projection,
@@
Thm. 4.14 not_orients_dup_rule, no_symbolic_variable_condition_orients_dup_step Symbolic schema Standard variable condition on the strict symbolic comparator.
Cor. 4.15 no_kbo_orients_ko7_rec_succ KO7 symbolic corollary KBO-style order satisfying the standard variable condition.
Cor. 4.16 no_affine_with_pump_orients_dup_step,
@@
Cor. 4.17 stepCtxFull_orientation_implies_root, no_stepCtxFull_orientation_affine_with_pump,
no_stepCtxFull_orientation_matrixFunctional_with_projected_affine_pump
@@
Thm. 4.18 no_additive_compositional_orients_rec_succ, no_compositional_orients_rec_succ_transparent_delta KO7 Tier 1: positive app weight. Tier 2: app-subterm axioms and delta-transparency at void.
@@
Thm. 4.21 dp_projection_orients_rec_succ, dp_projection_violates_sensitivity KO7 / DP None.
Prop. 4.22 counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order,
@@
no_affine_orients_alternating_dup2_composite_of_unbounded
Alternating SCC schema Additive branch: none. Affine branch: unbounded pump.
Thm. 4.24 step_preserves_payloadCount,
@@
no_affine_orients_synchronized_cycle_of_wrapper_dominance
@@
Cor. 4.25 no_global_orients_additive, no_global_orients_compositional_transparent_succ, no_global_orients_affine_of_unbounded Schema / global Affine branch inherits the unbounded-pump condition.
Prop. 4.26 W_orients_step, wf_StepRev_poly KO7 root Step None.
Prop. 4.27 wf_StepRev_mpo KO7 root Step Fixed signature, fixed precedence, fixed same-head clause.
Prop. 4.28 mpoOrd_lt_mpoVeblenBound, wf_StepRev_mpo_below_veblen7 KO7 root Step Fixed signature; ordinal envelope below Ordinal.veblen70.
Prop. 4.29 not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec KO7 root Step Bad-precedence hypothesis.
Thm. 4.30 ko7_direct_escape_trichotomy_extended (preliminary: ko7_nat_direct_escape_trichotomy; projection extension: ko7_projection_escape_trichotomy) KO7 direct universe Successful orienter must violate wrapper sensitivity, violate transparency, or leave the twelve explicit representable families.
Prop. 4.31 not_localJoinStep_eqW_refl KO7 root Step None.
@@
Prop. 5.13 safestep_length_bounded_by_size SafeStepCtx Size-indexed tower-exponential bound via ctxFuel.
Prop. 5.14 fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope,
@@
SafeStepCtx Explicit finite fast-growing-family envelope majorizing the same contextual length bound.

### 08 Apr 2026 v8

@@ 8 Related work
Table 11: Position of the present artifact relative to nearby formal developments.
Artifact Assistant Primary focus Fixed-system impossibility theorems? External certificate checking?
This artifact Lean 4 Fixed-system barrier theorems, escape classification, certified witnesses, and KO7 certification Yes Bridge / replay layer only
AFP undecidability formalization [28] Isabelle/HOL Undecidability of orientation problems across all inputs No No
IsaFoR/CeTA [25] Isabelle/HOL Certification and checking of positive TRS proofs No Yes
CoLoR / Coccinelle-CiME [5, 8] Coq Certified checking / soundness for positive rewriting arguments No Yes
Lean Metatheory [23] Lean 4 Generic rewriting, confluence, and normalization infrastructure No No
@@
Proof-assistant termination checkers.
The same obstruction is also relevant to proof assistants, even though the artifact does not formalize Lean’s termination_by checker, Agda’s Foetus checker, or Coq’s guard condition as theorem objects. The textbook duplicating rule $f(x,s(y))\to g(x,f(x,y))$, now packaged in Meta/TextbookDupInstance.lean, gives a concrete recursive pattern whose direct additive, transparent-compositional, and affine orientation already fails by theorem. So any termination checker restricted to direct structural or lexicographic aggregation of recursive arguments needs an escape mechanism analogous to argument projection or a stronger global witness: otherwise the duplicated payload remains on the wrong side of the boundary for exactly the same reason it defeats the direct measure classes formalized here.
@@ 9 Conclusion
We formalized an orientation boundary for first-order rewrite systems with step-duplicating recursors. Twelve direct measure classes in the barrier package are excluded by schema-level impossibility theorems (Theorems 4.1 to 4.14); the escape side is classified by the DP clause, the explicit DP base-order boundary, transparency essentiality, the generalized bounded-polynomial escape condition, the escape trichotomy, and the two SCC synchronization theorems; and the concrete witness calculus KO7 carries a guarded certification chain together with full-step polynomial and MPO witnesses.
The barrier stack separates baseline additive/compositional failure, widened scalar families, vector / pair projection layers, and the symbolic variable-condition side; the escape results track that organization closely. The paper therefore marks a boundary for explicit direct families, not a claim that KO7 defeats termination methods wholesale. Dependency pairs, LPO, the nonlinear witness, and the KO7-specialized MPO remain on the successful side of that line, and the repository states those imports at the theorem level.
The repository is a pinned Lean 4 snapshot for the explicit first-order setting treated here. It contains the schema-level impossibility theorems, the escape classification, the guarded certification chain, the full-step termination witnesses, and the external TRS validation trail.
@@ A.1 Lean kernel location
Path convention: manuscript references of the form Meta/... name Lean module paths in the OperatorKO7.Meta.* namespace; the corresponding repository-root files live under OperatorKO7/Meta/...lean. Literal filesystem paths are written repo-root-relative.
@@ A.2 Module map
Table LABEL:tab:module-map lists the theorem-bearing kernel and meta modules in the active artifact. Lightweight test files under OperatorKO7/Test/ are intentionally omitted.
@@
Meta/ContextClosedBarrier Root-to-context barrier survival corollaries for StepCtxFull (stepCtxFull_orientation_implies_root, representative entries no_stepCtxFull_orientation_affine_with_pump,
no_stepCtxFull_orientation_matrixFunctional_with_projected_affine_pump; see the module for the complete no_stepCtxFull_orientation_* family list)
@@
Meta/SafeStep_Complexity_Ordinal Legacy ctxFuel-based SafeStepCtx derivation-length bound; size-indexed tower-exponential corollary (safestep_length_bounded_by_size)
Meta/SafeStep_Complexity_FastGrowing Explicit finite fast-growing envelope majorizing the same coarse SafeStepCtx tower bound (fgOmegaEnvelope, safestep_length_bounded_by_fgOmegaEnvelope)
Meta/ContextualCopyBudget Positive position-aware contextual copy-budget analysis and strict descent proof for ctxDupPotential (ctxDupPotential_decreases_ctx, safeStepCtx_length_le_two_pow_double_termSize)
Meta/ContextualCopyBudget_NoGo Auxiliary contextual coordinates, obstruction witnesses, and the no-go theorem for the monotone arithmetic candidate class (coordExpr_payloadFamily_nondec, not_coordExprStackCertifies_payloadFamily)
Meta/SafeStepCtx_Complexity_Exponential Public single-exponential SafeStepCtx envelope refining the old tower bound (contextualExpBound, contextualExpBound_le_complexity_bound, safeStepCtx_length_le_contextualExpBound)
Meta/SafeStepCtx_Complexity_Cichon Finite-note cichon wrapper around the sharpened contextual exponential theorem (ctxExpNote, ctxExpCichonBound, safeStepCtx_length_le_ctxExpCichonBound)
Meta/SafeStepCtx_Complexity_LowerBound Explicit contextual lower family with exact lengths, linear source size, and exponential normal-form output size (ctxLowerFamily_pow_nf, safeStepCtx_has_singleExponential_lower_family,
safeStepCtx_has_singleExponential_output_family)
Meta/SafeStep_Complexity_MW_Root Notation bridge from the calibrated root measure to NONote, together with the genuine root cichon theorem (mwRootNote, safeStepPow_length_le_mwRootBound)
Meta/SafeStep_Complexity_MW_Ctx Conservative contextual cichon lift plus explicit obstruction to direct contextual reuse of the root note (safestepctx_length_le_mwCtxBound, not_all_safeStepCtx_rootNote_drop)
Meta/SafeStep_Complexity_MW_CtxExact Direct contextual exact controlled-descent package on a separate finite note family (ctxExactNote, safestepctx_length_le_ctxExactBound)
Meta/SafeRoot_Complexity Exact-path realization and explicit linear / tower / root-cichon envelopes for the certified root normalizer (normalizeSafeSteps_realized,
normalizeSafeSteps_le_linear_termSize, normalizeSafeSteps_le_complexity_bound, normalizeSafeSteps_le_mwRootBound)
Meta/Reachability_Complexity Explicit decision-cost envelope for safe-normal-form reachability, including a linear size upper bound and matching linear families (reachabilityDecisionCost_le_linear_termSize,
reachabilityDecision_has_linear_size_family)
Meta/SafeStepCtx_Confluence Discharged global contextual local-join obligation and unconditional SafeStepCtx confluence, retaining the Newman characterization as a corollary (localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll)
@@
Meta/Impossibility_Lemmas Failed-measure catalog; historical dead-end notes moved to Docs/Impossibility_DeadEnds.md
@@
Meta/BarrierWitness_Budgets Canonical witness-budget packaging for the extractor layer (WitnessStepShape, affine_with_pump_witness_shape,
quadratic_with_pump_witness_shape, max_with_pump_witness_shape)
@@
Meta/TextbookDupInstance Second named schema instance for the textbook rule $f(x,s(y))\to g(x,f(x,y))$, with direct additive/compositional/affine barrier corollaries and witness aliases (no_global_textbook_orientation_additive, textbook_additive_witness)
Meta/CompositionalMeasure_Impossibility KO7 instantiation; KO7 DP witness; KO7 affine barrier (Thms. 4.194.22;
@@
Meta/WPO_PolynomialBarrier WPO-facing direct polynomial-branch corollary lifting the bounded-polynomial barrier to direct orders sound for that algebra (no_wpoPolynomialDirect_orients_dup_step_of_unbounded,
no_global_step_orientation_wpoPolynomialDirect_of_succ_pump)
@@
Meta/MutualDuplication_CycleFlow Abstract delayed-duplication one-cycle metatheorems, packaging additive, affine, transparent-compositional, and scalar-projection contextual barriers from a certified cycle witness (no_global_orients_ctx_additive, no_global_orients_ctx_of_scalar_projection_affine_of_unbounded)
Meta/MutualDuplication_KNode Finite cyclic ‘k+1‘-node delayed-duplication generalization with exact cycle realization and additive / affine minimal-context barriers (cycle_realized, no_global_orients_ctx_affine_of_unbounded)
Meta/MutualDuplication_KNode_Abstract Bridge module showing that the finite delayed-duplication ‘k+1‘-node SCC factors through the abstract delayed-cycle layer, including transparent-compositional and projected matrix-style corollaries (no_global_orients_ctx_compositional_transparent_via_cycleFlow, no_global_orients_ctx_matrixFunctional_of_projected_unbounded_via_cycleFlow)
Meta/GraphPathExtraction Generic extraction of explicit edge paths from nonempty transitive-closure proofs, used to turn closed-path and round-trip SCC witnesses into concrete cycle data automatically (EdgePath, ofTransGen, ofRoundTrip)
Meta/FiniteGraphReachability Bounded successor-closure reachability for finite decidable directed graphs, proving that Fintype.card rounds capture reflexive transitive closure and supporting finite-round-trip SCC wrappers without explicit path proofs (Reachable, reachable_iff_reflTransGen, transGen_of_reachable_of_ne)
Meta/FiniteGraphSCC Finite-graph SCC witness packaging and search layer: define the finite pair-search surface and prove its equivalence with nontrivial SCC existence, then recover a concrete distinct pair together with both reachability directions internally (sccPairs, HasNontrivialSCC, witnessPair, reachable_witnessSrc_witnessDst)
Meta/MutualDuplication_GraphCycle Raw-graph delayed-duplication metatheorems for arbitrary node / edge systems equipped with a certified closed cycle witness, packaging additive, affine, transparent-compositional, and scalar-projection contextual barriers (CyclePath, no_global_orients_ctx_of_scalar_projection_affine_of_unbounded)
@@
Meta/MutualDuplication_PayloadFlow Abstract payload-flow metatheorems for one-cycle synchronized exposure, packaging additive and affine global contextual barriers independently of the concrete packet syntax (no_global_orients_ctx_additive, no_global_orients_ctx_affine_of_wrapper_dominance)
Meta/MutualDuplication_Preserving_KNode Finite cyclic synchronized-packet preserving generalization with exact payload preservation, one-cycle exposure, additive failure, and conditional affine wrapper-dominance failure (synchronized_cycle_realized, no_global_orients_ctx_affine_of_wrapper_dominance)
Meta/MutualDuplication_Preserving_Abstract Bridge module showing that the finite synchronized-packet SCC factors through the abstract payload-flow layer, including additive, affine, packet-transparent, and projected packet-transparent corollaries (no_global_orients_ctx_additive_via_payloadFlow, no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent_via_payloadFlow)
Meta/MutualDuplication_Preserving_Transparent Packet-transparent compositional and scalar-projection extension of the finite synchronized-packet preserving SCC barrier (no_global_orients_ctx_packetTransparent, no_global_orients_ctx_matrixFunctional_of_projected_packetTransparent)
Meta/MutualDuplication_PacketGraph Raw-graph synchronized-packet metatheorems for arbitrary node / edge systems equipped with a certified closed cycle witness, together with closed-path / round-trip extraction wrappers, packaging additive, affine, transparent, and scalar-projection contextual barriers (CyclePath, no_global_orients_ctx_of_scalar_projection_transparent_of_roundTrip)
Meta/MutualDuplication_RelationalGraph Relation-level construction layer building the delayed or preserving raw-graph SCC system automatically from a node type, edge relation, shared constructor interface, and local edge-realization theorem, then reexporting the round-trip, finite-round-trip, finite-SCC, and finite-search-result contextual barrier wrappers (no_global_orients_ctx_additive_of_hasNontrivialSCC, search-result wrappers for successful finite-SCC discovery)
Meta/MutualDuplication_CallGraph Call-graph construction layer building the delayed or preserving SCC system automatically from finite extracted ‘nodeKey‘ / ‘succKeys‘ data on a node type, then reexporting the same finite-search-result contextual barrier wrappers (no_global_orients_ctx_additive_of_hasNontrivialSCC, call-graph wrappers for successful finite-SCC discovery)
Meta/MutualDuplication_ExtractedCallGraph Array-backed extracted-data construction layer building the delayed or preserving call graph and SCC barrier interface automatically from raw extracted node data indexed by Finnodes.size (finite SCC search on the extracted graph, then additive / affine / transparent wrapper export)
@@
Meta/KBO_Impossible Paper-facing renaming layer over the symbolic variable-condition barrier in Meta/SymbolicComparatorBarrier.lean: re-exports the abstract obstruction under KBO-facing names and adds one genuinely new concrete-trace bridge theorem for rec_succ. The substantive proof lives in the symbolic barrier file, not here (no_kbo_orients_ko7_rec_succ, no_kbo_orients_ko7_rec_succ_trace)
@@
Meta/ManySortedBarrierSurvival Specialized many-sorted first-order repackaging of the typed recursor fragment; additive and affine barrier survival carry over unchanged (no_additive_orients_manySorted_recSucc,
no_affine_orients_manySorted_recSucc_of_stepPump)
Meta/DependencyPairs_Fragment Reusable DP fragment: rank-based pair problems and SCC-path descent (wfRev in namespace DPProjection, target_lt_source in namespace SCCCycle)
Meta/DependencyPairs_FiniteGraph Finite dependency-pair graph interface reusing the finite-SCC search layer and repackaging a discovered SCC as the fragment’s standard cycle witness (FiniteDPGraph, HasNontrivialSCC, toSCCCycle)
Meta/DependencyPairs_CallGraph Finite dependency-pair call-graph interface building the pair relation automatically from extracted ‘nodeKey‘ / ‘succKeys‘ data and reusing the same finite-SCC search / cycle packaging surface (FiniteCallGraph, HasNontrivialSCC, toSCCCycle)
Meta/DependencyPairs_ExtractedCallGraph Array-backed extracted dependency-pair call-graph interface generating the node type as Finnodes.size from raw extracted node data before reusing the same finite-SCC search / contradiction surface (FiniteExtractedCallGraph, ofArrayMap, HasNontrivialSCC)
Meta/DependencyPairs_TPDBExtraction Concrete TPDB-side dependency-pair extraction layer computing defined heads and right-hand-side call heads from a finite rule list, then building the array-backed extracted call graph automatically (tpdbDefinedHeads, extractTpdbNodes, tpdbExtractedCallGraph)
Meta/DependencyPairs_FirstOrderExtraction Generic first-order dependency-pair extraction layer over arbitrary symbols and variables, computing defined heads, extracted call-head nodes, and the array-backed call graph for any finite first-order TRS (FOTerm, FORule, foDefinedHeads, foExtractedCallGraph)
Meta/DependencyPairs_FirstOrderFrontend Rule-record frontend for the generic first-order extractor, taking a finite rule array plus lhs / rhs extractors directly to the extracted call graph and the same SCC-search / contradiction surface (extractRuleFrontendNodes, foExtractedCallGraphOf, HasNontrivialSCCOf)
Meta/DependencyPairs_FirstOrderEngine Consolidated engine-level frontend packaging an internal first-order TRS / DP engine as one object and reexporting the same extracted-call-graph and SCC-search surface; the same module now also contains the typeclass-based internal-engine view and the raw-term bridge for engines with their own syntax (FiniteFirstOrderEngine, HasFiniteFirstOrderView, HasFirstOrderTermView, HasFiniteRawFirstOrderView, extractedCallGraph)
Meta/DependencyPairs_HeadView Minimal head/call-head bridge for internal engines that only expose root-head and recursive-call-head data, rederiving the extracted call graph without a full term conversion, exposing a typeclass-level head-view engine surface, and adding explicit adapters from the first-order and raw-term engine layers (HasCallHeadView, FiniteHeadRuleEngine, HasFiniteHeadRuleView, finiteHeadRuleViewOfRaw, extractedCallGraph)
Meta/DependencyPairs_FiniteCarrierView Consolidated finite-carrier bridge for internal engines whose rules already form a finite type, rederiving the canonical first-order, raw-term, head-view, and extracted-data surfaces from one module (FiniteCarrierFirstOrderEngine, HasFiniteCarrierFirstOrderView, HasFiniteCarrierRawFirstOrderView, FiniteCarrierHeadEngine, HasFiniteCarrierHeadView, FiniteCarrierExtractedEngine, HasFiniteCarrierExtractedView, finiteCarrierExtractedViewOfHeadCarrier)
Meta/DependencyPairs_KernelFirstOrder Internal KO7-symbol frontend for the generic first-order extractor, packaging the eight full KO7 rules over an internal symbol type, then exposing them through HasFiniteRawFirstOrderView, deriving the smaller head-view engine through the generic adapter layer, and feeding the same system through the consolidated finite-carrier stack before rederiving the extracted-node / call-graph surfaces without TPDB syntax (Symbol, EngineTag, RuleId, ko7Engine, ko7HeadEngine, ko7CarrierFirstOrderEngine, ko7CarrierHeadOnlyEngine, ko7CarrierExtractedGraph, ko7FullStepExtractedCallGraph)
@@
Meta/Mu3c_Image_LowerBound Explicit realized lower-bound families for the full triple-lex image in both $\delta$-blocks (mu3c_lower_family_block0,
mu3c_lower_family_block1)
@@
Meta/PolyInterpretation_FullStep Nonlinear polynomial $W$; axiom-violation witnesses (Rem. 4.21;
@@ A.3 Relation/property status matrix
Table 13: Certified status matrix for the main relations and transformed systems.
@@
Step Root termination: yes Root confluence: no (witness); normalizer: n/a nonlinear polynomial and MPO witnesses; MPO ranking bounded above by $\varphi_{7}(0)$ wf_StepRev_poly, wf_StepRev_mpo, wf_StepRev_mpo_below_veblen7, not_localJoinStep_eqW_refl
SafeStep Strong normalization: yes Root confluence + certified normalizer + reachability to safe normal-form targets triple-lex ordinal upper-bound calibration, explicit two-block lower families in the realized mu3c image, explicit linear root-normalizer lower/upper envelope, plus the root cichon corollary wf_SafeStepRev_c, confluentSafe, normalizeSafe_sound, reachability_decidable, mu3c_lower_family_block0, mu3c_lower_family_block1, normalizeSafeSteps_le_linear_termSize, normalizeSafeSteps_le_mwRootBound, safeMeasure_below_epsilon0
SafeStepCtx Partial-context SN: yes Confluence: yes; Newman characterization retained single-exponential structural bound with matching explicit lower family, retained tower / fast-growing corollaries, conservative calibrated cichon lift, and direct exact contextual cichon lift on a separate note family wf_SafeStepCtxRev, localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, safeStepCtx_length_le_contextualExpBound, safestep_length_bounded_by_size, safestep_length_bounded_by_fgOmegaEnvelope, safestepctx_length_le_mwCtxBound, safestepctx_length_le_ctxExactBound
@@ A.4 Claim-to-code index
Tables LABEL:tab:claim-code-barrier and LABEL:tab:claim-code-certification index the numbered paper results by Lean identifier, abstraction level, and side conditions. This complements the module map: the tables below are organized by paper statement rather than by file. When a paper result is backed by a bridge theorem together with a family of repetitive class-specific corollaries, the tables list the bridge and the primary anchors rather than exhaustively enumerating every specialization.
@@
Cor. 4.7 no_wpoPolynomialDirect_orients_dup_step_of_unbounded, no_global_step_orientation_wpoPolynomialDirect_of_succ_pump,
wpoPolynomialDirect_escape_requires_failure_of_base_dominance
Schema / KO7 interface corollary A direct order whose strict comparison is sound for the bounded polynomial algebra of Thm. 4.6.
Thm. 4.8 no_max_orients_dup_step_of_unbounded, no_global_step_orientation_max_of_succ_pump Schema / KO7 Max-plus constructor-local interface with unbounded range, or positive successor / left-wrapper drift.
Cor. 4.9 no_arctic_primary_orients_dup_step_of_unbounded, no_global_step_orientation_arctic_primary_of_unbounded Schema / KO7 Distinguished finite arctic primary projection satisfying the max-plus interface.
Thm. 4.10 no_matrixD_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with one tracked affine coordinate and growth in that coordinate.
Thm. 4.11 no_matrix2_lex_orients_dup_step_of_unbounded_primary Schema / KO7 Lexicographic order by the first component; unbounded or pumped primary coordinate.
Thm. 4.12 no_matrixMix2_orients_dup_step_of_sum_pump Schema / KO7 Balanced mixed $2\times 2$ constructor maps and an unbounded aggregate-sum pump.
Thm. 4.13 no_matrixFunctional_orients_dup_step_of_componentwise_pump Schema / KO7 Strict componentwise order with a fixed nonzero natural-weighted scalar projection satisfying the affine interface.
Thm. 4.14 no_orients_dup_step_of_scalar_projection,
@@
Thm. 4.15 not_orients_dup_rule, no_symbolic_variable_condition_orients_dup_step Symbolic schema Standard variable condition on the strict symbolic comparator.
Cor. 4.16 no_kbo_orients_ko7_rec_succ, no_kbo_orients_ko7_rec_succ_trace (paper-facing renaming layer over Meta/SymbolicComparatorBarrier.lean; substantive proof: not_orients_dup_rule) KO7 symbolic corollary KBO-style order satisfying the standard variable condition.
Cor. 4.17 no_affine_with_pump_orients_dup_step,
@@
Cor. 4.18 stepCtxFull_orientation_implies_root, no_stepCtxFull_orientation_additive_compositional,
no_stepCtxFull_orientation_compositional_transparent_delta,
no_stepCtxFull_orientation_affine_with_pump, no_stepCtxFull_orientation_quadratic_with_pump,
no_stepCtxFull_orientation_cross_quadratic_with_pump, no_stepCtxFull_orientation_multilinear_with_pump,
no_stepCtxFull_orientation_max_with_pump, no_stepCtxFull_orientation_matrix2_with_primary_pump,
no_stepCtxFull_orientation_matrix2_lex_with_primary_pump,
no_stepCtxFull_orientation_matrixFunctional_with_projected_affine_pump,
no_stepCtxFull_orientation_matrixMix2_with_sum_pump
@@
Thm. 4.19 no_additive_compositional_orients_rec_succ, no_compositional_orients_rec_succ_transparent_delta KO7 Tier 1: positive app weight. Tier 2: app-subterm axioms and delta-transparency at void.
@@
Thm. 4.22 dp_projection_orients_rec_succ, dp_projection_violates_sensitivity KO7 / DP None.
Prop. 4.23 counterHeight_orients_dpPair, extracted_dp_problem_has_linear_base_order,
@@
no_affine_orients_alternating_dup2_composite_of_unbounded,
cycle_realized,
no_global_orients_ctx_affine_of_unbounded
Alternating SCC schema Additive branch: positive wrap weight inherited from Thm. 4.1. Affine branch: positive wrap weight + unbounded pump.
Thm. 4.25 step_preserves_payloadCount,
@@
no_affine_orients_synchronized_cycle_of_wrapper_dominance,
no_global_orients_ctx_affine_of_wrapper_dominance
@@
Cor. 4.26 no_global_orients_additive, no_global_orients_compositional_transparent_succ, no_global_orients_affine_of_unbounded Schema / global Affine branch inherits the unbounded-pump condition.
Prop. 4.27 W_orients_step, wf_StepRev_poly KO7 root Step None.
[…]
Prop. 4.29 mpoOrd_lt_mpoVeblenBound, wf_StepRev_mpo_below_veblen7 KO7 root Step Fixed signature; ordinal envelope below Ordinal.veblen70.
Prop. 4.30 not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec KO7 root Step Bad-precedence hypothesis.
Thm. 4.31 ko7_direct_escape_trichotomy_extended (preliminary: ko7_nat_direct_escape_trichotomy; projection extension: ko7_projection_escape_trichotomy) KO7 direct universe Successful orienter must violate wrapper sensitivity, violate transparency, or leave the twelve explicit representable families.
Prop. 4.32 not_localJoinStep_eqW_refl KO7 root Step None.
@@
Prop. 5.14 ctxDupPotential, ctxDupPotential_decreases_ctx,
copyBudget_le_termSize, ctxDupPotential_add_one_le_two_pow_double_termSize,
safeStepCtx_length_le_two_pow_double_termSize, contextualExpBound,
contextualExpBound_le_complexity_bound, safeStepCtx_length_le_contextualExpBound
SafeStepCtx Single-exponential structural bound via the position-aware multiplicity potential ctxDupPotential; the older ctxFuel/towerBound estimate survives only as a coarser corollary.
Prop. 5.15 ctxLowerFamily, ctxLowerNF, ctxLowerLen,
termSize_ctxLowerFamily, termSize_ctxLowerNF, ctxLowerFamily_pow_nf,
two_pow_le_ctxLowerLen, safeStepCtx_has_singleExponential_lower_family,
safeStepCtx_has_singleExponential_output_family
SafeStepCtx Explicit linear-size contextual lower family with exponentially long exact chains and exponential normal-form output size, matching the new guarded-context upper theorem up to constants in the exponent and giving a direct duplication-growth witness.
Prop. 5.16 fgOmegaEnvelope, towerBound_le_fgOmegaEnvelope,
@@
SafeStepCtx Explicit finite fast-growing-family envelope majorizing the older coarse contextual tower bound.
Prop. 5.17 mwCtxNote, mwCtxBound, mwCtxNote_lt_opow_omega_mul_two,
ctxFuel_le_mwCtxBound, safestepctx_length_le_mwCtxBound,
not_all_safeStepCtx_rootNote_drop
SafeStepCtx Conservative ctxFuel-to-lex3Note lift; direct contextual reuse of the root note is explicitly obstructed.
Prop. 5.18 ctxExactNote, ctxExactBound, ctxExactNote_lt_omega,
safeStepCtx_exact_step, safeStepCtxPow_exact_drop,
safestepctx_length_le_ctxExactBound
SafeStepCtx Direct exact contextual fundamental-sequence extraction on a separate finite note family built from ctxFuel.